OpenStudy (anonymous):

PLEASE HELP!!! GIVING MEDALS!!! What is the length of the midsegment of the trapezoid made by the vertices A(0,5)B(3,3)C(5,-2) and D(-1,2). Show equations and all work that lead up to your answer.

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

midpoint formula: \[(\frac{ x _{1}+x _{2} }{ 2 }, \frac{ y _{1}+y _{2} }{ 2 })\] try between points (0,5) and (-1, 2) then (3, 3) and (5,-2) then distance formula is: \[\sqrt{(x _{2}-x _{1})^{2} + (y_{2}-y _{1})^{2}}\] It's past my bed time >,< I'm gonna crash, good night.

OpenStudy (anonymous):

Just one more question! PLEASE

OpenStudy (anonymous):